OpenStudy (itsbribro):

Which mathematical property is demonstrated? 72+(28+93)=(72+28)+93 A. communicative property of addition B. associative property of addition

OpenStudy (anonymous):

This is the associative property of addition. The communicative property is when you change the order of the numbers themselves (ex: 10 + 4 = 4 + 10). The associative property changes the parentheses.

OpenStudy (itsbribro):

can someone tell me what every property is and not just the ones that are listed here.

OpenStudy (mikeymaximum):

The answer is B. associative property of addition

OpenStudy (mikeymaximum):

The Commutative Property of Addition. The Commutative Property of Multiplication. The Associative Property of Addition. The Associative Property of Multiplication. The Distributive Property. The Additive Identity Property. The Additive Inverse Property. The Multiplicative Identity Property

OpenStudy (itsbribro):

okay thank you @TheChancellor and @MikeyMaximum

OpenStudy (itsbribro):

yes can you give me an example of each?

OpenStudy (mikeymaximum):

"The word "commutative" comes from "commute" or "move around", so the Commutative Property is the one that refers to moving stuff around. For addition, the rule is "a + b = b + a"; in numbers, this means 2 + 3 = 3 + 2. For multiplication, the rule is "ab = ba"; in numbers, this means 2×3 = 3×2."

OpenStudy (anonymous):

The other properties (at this level of math) are the identity property which states any number added to 0 will be that number (ex: 0 + x = x) and the distributive property (ex: a * (b + c) = (a * b) + (a * c)... good luck!

OpenStudy (mikeymaximum):

"Since all they did was regroup things, this is true by the Associative Property. The word "commutative" comes from "commute" or "move around", so the Commutative Property is the one that refers to moving stuff around. For addition, the rule is "a + b = b + a"; in numbers, this means 2 + 3 = 3 + 2."

OpenStudy (mikeymaximum):

"The Distributive Property is an algebra property which is used to multiply a single term and two or more terms inside a set of parentheses."

OpenStudy (itsbribro):

thank you both so much <3 :3

OpenStudy (itsbribro):

can you help me with a couple more questions @MikeyMaximum

OpenStudy (mikeymaximum):

I can try

OpenStudy (itsbribro):

just 2 more

OpenStudy (mikeymaximum):

But do you still need explanations?

OpenStudy (itsbribro):

no thank you :)

OpenStudy (mikeymaximum):

OK

OpenStudy (mikeymaximum):

Can you kinda like Close this Q so I can get a medal or... Is that too much at the moment?

OpenStudy (itsbribro):

Which mathematical property is demonstrated? If x=-3 and -3=z, then x=z A. symmetric property of equality B. transitive property of equality